# Continual Low-Rank Scaled Dot-product Attention

### v2

@@ V-C Audio Classification experiments
In our experiments, we used the GTZAN Music Genre Classification dataset [43], which contains 100 music tracks for each of the 10 different music genres. Each track has a length of 30 seconds. Following the same architecture as previous works [28], we generate the Mel Spectrogram of each track. This spectrogram is an image where the width and height are the temporal and feature dimensions, respectively. Then, we divide it into one-second clips with a stride of 0.25 seconds, resulting in 120 clips. These are then introduced to a VGGish network [44, 45] pre-trained on the AudioSet dataset [45], resulting in a sequence of 120 tokens ready to be used as the input to a Transformer model.
@@
We performed the experiments with five different data splits between the training, validation and test sets and five different random weight initializations, resulting in 25 runs. For each of the competing methods, we evaluated two models with one and two Transformer layers, respectively, each having 16 heads in the SDA module and token dimensionality of $d=192$. Table II provides the results obtained by using models having the competing SDA formulations. In the Table, FLOPs are counted in millions, the last column references the relative amount of FLOPs of each SDA in relation to the original SDA with the same number of transformer layers, and all Nyström-based models in that table use $m=4$ landmarks. The results in accuracy are similar in general terms. The models with the original SDA and Continual SDA formulations achieve a slightly higher accuracy compared to the models having a Nyström-based SDA formulation. Using two Transformer layers results in a slight increase in accuracy. As for the number of computations, the continual models are significantly more efficient, with the model using the proposed SDA formulation with fixed landmarks having the lowest computational cost.
@@ V-D Online Action Detection experiments
Table III provides the results obtained by using the Segment Network model pre-trained on the Kinetics-400 [51] and the ActivityNet [50] datasets, respectively. As can be seen in these Tables, models using Nyström-based SDA formulations achieve lower mAP values compared to the models using the original SDA and the Continual SDA formulations, in exchange for a reduced computational cost. The use of fixed landmarks models translates in another reduction in mAP. In terms of cmAP, we can see that the performance of all models is very similar. No significant difference in performance is observed between the models using continual SDA formulations and their non-continual counterparts.

### v3

@@ V-C Audio Classification experiments
In our experiments, we used the GTZAN Music Genre Classification dataset [44], which contains 100 music tracks for each of the 10 different music genres. Each track has a length of 30 seconds. Following the same architecture as previous works [28], we generate the Mel Spectrogram of each track. This spectrogram is an image where the width and height are the temporal and feature dimensions, respectively. Then, we divide it into one-second clips with a stride of 0.25 seconds, resulting in 120 clips. These are then introduced to a VGGish network [45, 46] pre-trained on the AudioSet dataset [46], resulting in a sequence length ($n$) of 120 tokens ready to be used as the input to a Transformer model.
@@
We performed the experiments with five different data splits between the training, validation and test sets and five different random weight initializations, resulting in 25 runs. For each of the competing methods, we evaluated two models with one and two Transformer layers, respectively, each having 16 heads in the SDA module and token dimensionality of $d=192$. Table III provides the results obtained by using models having the competing SDA formulations. In the table, FLOPs are counted in millions, the last column references the relative amount of FLOPs of each SDA in relation to the original SDA with the same number of Transformer layers, and all Nyström-based models in that table use $m=4$ landmarks. The results in accuracy are similar in general terms. The models with the original SDA and Continual SDA formulations achieve a slightly higher accuracy compared to the models having a Nyström-based SDA formulation. Using two Transformer layers results in a slight increase in accuracy. As for the number of computations, the continual models are significantly more efficient, with the model using the proposed SDA formulation with fixed landmarks having the lowest computational cost.
@@ V-D Online Action Detection experiments
Table IV provides the results obtained by using the Segment Network model pre-trained on the Kinetics-400 [52] and the ActivityNet [51] datasets, respectively. As can be seen in these tables, models using Nyström-based SDA formulations achieve lower mAP values compared to the models using the original SDA and the Continual SDA formulations, in exchange for a reduced computational cost. The use of fixed landmarks models translates in another reduction in mAP. In terms of cmAP, we can see that the performance of all models is very similar. No significant difference in performance is observed between the models using continual SDA formulations and their non-continual counterparts.
@@ VI Conclusion
-A Computational costs calculation
This section shows the exact theoretical computational costs of the different attention modules. In particular, we show the cost of performing a single update during inference time on each of the variants of the Transformers discussed in this work. The real cost of these may be different based on the specific implementation and low-level operations.
In the following sections, $n$ makes reference to the sequence length, $d$ makes reference to the number of features and $m$ to the number of landmarks for the corresponding Nyström approximations.
-A1 Base Transformer
The computational cost is $2n^{2}d+n^{2}+nd+n$:
  • $nd$ from normalizing the matrix $Q$
  • $n^{2}d$ to perform the multiplication $\frac{QK^{T}}{\sqrt{d}}$
  • $n^{2}+n$ to perform the softmax activation
  • $n^{2}d$ to multiply the resulting matrix $\text{s}\left(\frac{QK^{T}}{\sqrt{d}}\right)$ against $V$
-A2 Nyströmformer
The computational cost is $4ndm+2nd+n+nm^{2}+2nm+dm^{2}+24m^{3}+22m^{2}+2m$:
  • $2nd$ from computing the landmark matrices $\tilde{Q}$ and $\tilde{K}$
  • $2ndm+2nm+n+m$ from computing the matrices $\text{s}\left(\frac{Q\tilde{K}^{T}}{\sqrt{d}}\right)$ and $\text{s}\left(\frac{\tilde{Q}K^{T}}{\sqrt{d}}\right)$
  • $dm^{2}+m^{2}+m$ from computing the matrix $\text{s}\left(\frac{\tilde{Q}\tilde{K}^{T}}{\sqrt{d}}\right)$
  • $24m^{3}+21m^{2}$ from computing the Moore-Penrose pseudo-inverse using the iterative algorithm [38]
  • $nm^{2}+2ndm$ from the multiplication of the 3 low-rank matrices
-A3 Nyströmformer with fixed landmarks
The computational cost is $4ndm+nm^{2}+2nm+n+m$:
  • $2ndm+2nm+n+m$ from computing the matrices $\text{s}\left(\frac{Q\tilde{K}^{T}}{\sqrt{d}}\right)$ and $\text{s}\left(\frac{\tilde{Q}K^{T}}{\sqrt{d}}\right)$
  • $nm^{2}+2ndm$ from the multiplication of the 3 low-rank matrices
-A4 Continual Retroactive Transformers
The computational cost is $27nd+4n-2d-$ 2:
  • $nd+n-1$ from computing $\rho\left(Q_{\text{mem}},k_{\text{old}}\right)$ and $\rho\left(Q_{\text{mem}},k_{\text{new}}\right)$
  • $nd+n$ from computing the matrix $\rho\left(q_{\text{new}},K\right)$
  • $n$ from computing the vector $\phi(A)_{\text{new}}$
  • $nd$ from computing the vector $AV_{\text{new}}$
  • $n-1$ from updating $d_{\text{mem}}$
  • $3nd-2d$ from updating $AV_{\text{mem}}$
  • $nd$ from computing $\phi(A)^{-1}\odot AV$
-A5 Continual Single Output Transformer
The computational cost is $3nd+2n$:
  • $nd+n$ from computing the matrix $a=\rho\left(q_{\text{new}},K\right)$
  • $nd$ from computing the matrix $aV$
  • $n$ from computing the vector $\phi(a)$
  • $nd$ from computing $aV\phi(a)^{-1}$
-A6 Continual Retroactive Nyströmformers with Fixed Landmarks
The computational cost is $ndm+6dm+m^{2}+6m$:
  • $dm+3m$ from computing $(\beta_{\phi})_{\text{new}}$
  • $m^{2}$ from computing $(\beta_{\phi})_{\text{new}}(\Gamma_{\phi})^{\dagger}$
  • $2dm+2m$ from computing $\rho(\tilde{Q},k_{\text{old}})$ and $\rho(\tilde{Q},k_{\text{new}})$
  • $m$ from updating $\phi(\Delta)^{-1}$
  • $3dm$ from updating $\Delta V$
  • $ndm$ to complete the attention computation
-A7 Continual Single Output Nyströmformers with Fixed Landmarks
The computational cost is $7dm+m^{2}+6m$:
  • $dm+3m$ from computing $(\beta_{\phi})_{\text{new}}$
  • $m^{2}$ from computing $(\beta_{\phi})_{\text{new}}(\Gamma_{\phi})^{\dagger}$
  • $2dm+2m$ from computing $\rho(\tilde{Q},k_{\text{old}})$ and $\rho(\tilde{Q},k_{\text{new}})$
  • $m$ from updating $\phi(\Delta)^{-1}$
  • $3dm$ from updating $\Delta V$
  • $dm$ to complete the attention computation
-A8 Continual Retroactive Nyströmformers with Continual Landmarks
The computational cost is $ndm+8nd+nm^{2}+nm+11n+15dm+2d+24m^{3}+22m^{2}+22m$:
  • $dm+m$ from computing $\beta_{\text{new}}$
  • $nd+n$ from computing $\rho(Q_{\text{mem}},\tilde{k}_{\text{new}}^{T})$
  • $2d+m$ from updating $\Gamma$
  • $nd+n+dm+m$ from updating $\Delta$
  • $2nd+4n+dm+2m$ from updating $\phi(B)$
  • $3dm+6m$ from updating $\phi(\Gamma)$
  • $nd+2n+2dm+4m$ from updating $\phi(\Delta)$
  • $24m^{3}+21m^{2}$ from computing the Moore-Penrose pseudo-inverse using the iterative algorithm [38]
  • $nm^{2}+nm+m^{2}$ from computing $B_{\phi}(\Gamma_{\phi})^{\dagger}$
  • $4dm+2m$ from computing $\rho(\tilde{Q},k_{\text{old}})v_{\text{old}}$ and $\rho(\tilde{Q},k_{\text{new}})v_{\text{new}}$
  • $nd+2n+2dm+4m$ from computing $\phi(\Delta)$
  • $2nd+n$ from computing $\Delta V$
  • $dm+m$ from computing $\phi(\Delta)^{-1}\odot(\Delta V)$
  • $ndm$ to complete the attention multiplication
-A9 Continual Single Output Nyströmformers with Continual Landmarks
The computational cost is $ndm+3nd+n+9dm+2d+24m^{3}+22m^{2}+13m$:
  • $24m^{3}+21m^{2}$ from computing the Moore-Penrose pseudo-inverse using the iterative algorithm [38]
  • $2d+m$ from updating $(\Gamma_{\phi})^{\dagger}$
  • $2dm+m^{2}+3m$ from computing the updated $(\beta_{\phi})_{\text{new}}(\Gamma_{\phi})^{\dagger}$
  • $3nd+n+7dm+9m$ from updating $\Delta V$
  • $ndm$ to complete the attention multiplication
-B Memory costs calculation
Computing the memory cost of the SDA module for Online Inference is not trivial. First, we need to distinguish the so-called valley cost, i.e., the minimum cost corresponding to the memory when no inference is performed and all the necessary matrices are stored, and the so-called peak cost, corresponding the maximum memory needed for performing calculations using intermediate matrices and storing the necessary matrices during an inference step.
Since the main objective in Continual Inference is to improve processing speed by reducing redundant computations, optimizations (such as caching a compute result that can be used later resulting to a small memory increase in exchange for a computational cost reduction) are applied. Moreover, since specific implementation choices and low-level operations may lead to differences in the actual memory allocations of different implementations, in the following we consider the cases corresponding to the minimum possible cost.
-B1 Base Transformer
The valley cost is $3(nd-1)$:
  • $3(nd-1)$ from storing the matrices $Q_{\text{mem}}$, $K_{\text{mem}}$ and $V_{\text{mem}}$.
The peak cost is $n^{2}+4nd+1$:
  • $3nd$ from storing the matrices $Q,K,V$
  • $n^{2}$ from getting the matrix $\frac{QK^{T}}{\sqrt{d}}$
  • $1$ from performing the softmax operation
  • $nd$ from storing the output matrix
-B2 Nyströmformer
The valley cost is $3(nd-1)$:
  • $3(nd-1)$ from storing the matrices $Q_{\text{mem}}$, $K_{\text{mem}}$ and $V_{\text{mem}}$.
The peak cost is $4nd+2nm+2dm+1+6m^{2}+m$:
  • $3nd$ from storing the matrices $Q,K,V$
  • $2dm$ from storing the matrices $\tilde{Q},\tilde{K}$
  • $nm+1$ from storing the matrix $\text{s}(\frac{Q\tilde{K}^{T}}{\sqrt{d}})$
  • $m^{2}$ from storing the matrix $\text{s}(\frac{\tilde{Q}\tilde{K}^{T}}{\sqrt{d}})$ (we reuse the +1 from the previous operation for the softmax).
  • $5m^{2}+m$ from computing the Moore-Penrose pseudo-inverse intermediate results
  • $nm$ from storing the matrix $\text{s}(\frac{\tilde{Q}K^{T}}{\sqrt{d}})$
  • $nd$ to compute the intermediate and final results.
-B3 Nyströmformer with fixed landmarks
The valley cost is $3(nd-1)+2dm+m^{2}$:
  • $3(nd-1)$ from storing the matrices $Q_{\text{mem}}$, $K_{\text{mem}}$ and $V_{\text{mem}}$
  • $2dm$ from storing the matrices $\tilde{Q},\tilde{K}$
  • $m^{2}$ from storing the $(\Gamma_{\phi})^{\dagger}$ matrix
The peak cost is $14nd+2nm+2dm+2m^{2}+$ 1:
  • $3nd$ from storing the matrices $Q,K,V$
  • $2dm$ from storing the matrices $\tilde{Q},\tilde{K}$
  • $m^{2}$ from storing the $(\Gamma_{\phi})^{\dagger}$ matrix
  • $nm+1$ from storing the matrix $\text{s}(\frac{Q\tilde{K}^{T}}{\sqrt{d}})$
  • $nm$ from storing the matrix $\text{s}(\frac{\tilde{Q}K^{T}}{\sqrt{d}})$
  • $nd$ to compute the intermediate and final results.
-B4 Continual Retroactive Transformers
The valley cost is $44nd+n-d-$ 4:
  • $3(nd-1)$ from storing the matrices $Q_{\text{mem}}$, $K_{\text{mem}}$ and $V_{\text{mem}}$
  • $n-1$ from storing the vector $\phi(\hat{A}_{\text{prev}})$
  • $nd-d$ from storing the matrix $\hat{AV}_{\text{prev}}$.
The peak cost is $5nd+2n$:
  • $3nd$ from storing the matrices $Q$, $K$ and $V$
  • $n$ from storing the vector $\phi(A)$
  • $nd$ from storing the matrix $AV$.
  • $n$ from storing the result of $\rho(Q_{\text{mem}},k_{\text{old}})$ (we can also use this to store $\rho(Q_{\text{mem}},k_{\text{new}})$ and $\rho(q_{\text{new}},K)$)
  • $nd$ to store the final result.
-B5 Continual Single Output Transformer
The valley cost is $2(nd-1)$:
  • $2(nd-1)$ from storing the matrices $K_{\text{mem}}$ and $V_{\text{mem}}$
The peak cost is $2nd+n+2d$:
  • $2nd$ from storing the matrices $K$ and $V$
  • $d$ for storing $q_{\text{new}}$
  • $n$ for storing $\rho(q_{\text{new}},K)$
  • $d$ to store the final result
-B6 Continual Retroactive Nyströmformers with Fixed Landmarks
The valley cost is $nm+3dm+m^{2}+m$:
  • $2dm$ from storing the matrices $\tilde{Q}$ and $\tilde{K}$
  • $nm-m$ from storing the matrix $(B_{\phi}(\Gamma_{\phi})^{\dagger})_{\text{mem}}$
  • $m^{2}$ from storing the matrix $(\Gamma_{\phi})^{\dagger}$.
  • $m$ from storing the cached vector $\phi(\Delta)^{-1}$
  • $dm$ from storing the matrix $\Delta V$
The peak cost is $nd+nm+3dm+m^{2}+2m$:
  • $2dm$ from storing the matrices $\tilde{Q}$ and $\tilde{K}$
  • $nm$ from storing the matrix $(B_{\phi}(\Gamma_{\phi})^{\dagger})$
  • $m^{2}$ from storing the matrix $(\Gamma_{\phi})^{\dagger}$
  • $m$ from storing the cached vector $\phi(\Delta)^{-1}$
  • $dm$ from storing the matrix $\Delta V$
  • $m$ to store $\rho(\tilde{Q},k_{\text{old}})$
  • $nd$ to store the final result
-B7 Continual Single Output Nyströmformers with Fixed Landmarks
The valley cost is $3dm+m^{2}+m$:
  • $2dm$ from storing the matrices $\tilde{Q}$ and $\tilde{K}$
  • $m^{2}$ from storing the matrix $(\Gamma_{\phi})^{\dagger}$
  • $m$ from storing the cached vector $\phi(\Delta)^{-1}$
  • $dm$ from storing the matrix $\Delta V$
The peak cost is $3dm+d+m^{2}+2m$:
  • $2dm$ from storing the matrices $\tilde{Q}$ and $\tilde{K}$
  • $m^{2}$ from storing the matrix $(\Gamma_{\phi})^{\dagger}$
  • $m$ from storing the cached vector $\phi(\Delta)^{-1}$
  • $dm$ from storing the matrix $\Delta V$
  • $m$ to store $\phi(\tilde{Q},k_{\text{old}})$ and $(\beta_{\phi}(\Gamma_{\phi})^{\dagger})_{\text{new}}$
  • $d$ to store the final result
-B8 Continual Retroactive Nyströmformers with Continual Landmarks
The valley cost is $63nd+2nm+4dm+2m^{2}-2m-$ 6:
  • $3(nd-1)$ from storing the matrices $Q_{\text{mem}}$, $K_{\text{mem}}$ and $V_{\text{mem}}$
  • $2dm$ from storing the matrices $\tilde{Q}$ and $\tilde{K}$
  • $nm-m$ from storing the matrix $(B_{\phi}(\Gamma_{\phi})^{\dagger})_{\text{mem}}$
  • $m^{2}$ from storing the matrix $(\Gamma_{\phi})^{\dagger}$
  • $m$ from storing the cached vector $\phi(\Delta)^{-1}$
  • $dm$ from storing the matrix $\Delta V$
  • $n-1+2(m-1)$ for storing $\phi(\hat{B})_{\text{prev}}$, $\phi(\hat{\Gamma})_{\text{prev}}$ and $\phi(\hat{\Delta})_{\text{prev}}$
  • $nm-n-m$ from storing $B_{\text{prev}}$
  • $m^{2}-2m$ from storing $\Gamma_{\text{prev}}$
  • $dm-m$ from storing $(\Delta V)_{\text{prev}}$
The peak cost is $4nd+3nm+2n+4dm+7m^{2}+4m$:
  • $3nd$ from storing the matrices $Q$, $K$ and $V$
  • $2dm$ from storing the matrices $\tilde{Q}$ and $\tilde{K}$
  • $nm$ from storing the matrix $B_{\phi}(\Gamma_{\phi})^{\dagger}$
  • $m^{2}$ from storing the matrix $(\Gamma_{\phi})^{\dagger}$
  • $m$ from storing the cached vector $\phi(\Delta)^{-1}$
  • $dm$ from storing the matrix $\Delta V$
  • $n+2m$ for storing $\phi(\hat{B})$, $\phi(\hat{\Gamma})$ and $\phi(\hat{\Delta})$
  • $nm$ from storing $B$
  • $m^{2}$ from storing $\Gamma$
  • $dm$ from storing $(\Delta V)$
  • $n+m$ from 2 auxiliary vectors during computation
  • $5m^{2}$ from the Moore-Penrose pseudo-inverse
  • $nm$ to store the intermediate result $\Delta_{\phi}V$
  • $nd$ to store the final computation
-B9 Continual Single Output Nyströmformers with Continual Landmarks
The valley cost is $22nd+4dm+2m^{2}-$ 2:
  • $2(nd-1)$ from storing the matrices $K_{\text{mem}}$ and $V_{\text{mem}}$
  • $2dm$ from storing the matrices $\tilde{Q}$ and $\tilde{K}$
  • $m^{2}$ from storing the matrix $(\Gamma_{\phi})^{\dagger}$
  • $m$ from storing the cached vector $\phi(\Delta)^{-1}$
  • $dm$ from storing the matrix $\Delta V$
  • $2(m-1)$ for storing $\phi(\hat{\Gamma})_{\text{prev}}$ and $\phi(\hat{\Delta})_{\text{prev}}$
  • $m^{2}-2m$ from storing $\Gamma_{\text{prev}}$
  • $dm-m$ from storing $(\Delta V)_{\text{prev}}$
The peak cost is $2nd+nm+n+4dm+d+7m^{2}+4m$:
  • $2nd$ from storing the matrices $K$ and $V$
  • $d$ from storing the vector $q_{\text{new}}$
  • $2dm$ from storing the matrices $\tilde{Q}$ and $\tilde{K}$
  • $m^{2}$ from storing the matrix $(\Gamma_{\phi})^{\dagger}$
  • $m$ from storing the cached vector $\phi(\Delta)^{-1}$
  • $dm$ from storing the matrix $\Delta V$
  • $2m$ for storing $\phi(\hat{\Gamma})$ and $\phi(\hat{\Delta})$
  • $m^{2}$ from storing $\Gamma$
  • $dm$ from storing $(\Delta V)$
  • $n+m$ from 2 auxiliary vectors during computation
  • $5m^{2}$ from the Moore-Penrose pseudo-inverse
  • $nm$ to store the intermediate result $\Delta_{\phi}V$